A natural gas power plant uses 505Mg of natural gas per day (volume converted to weight via density ). The energy value of the natural gas is 37.3 MJ/kg. The plant produces 1.81 *10^6kWh of electricity each day

a.) what is the efficiency of the plant

b.) how much energy is lost please include all base equations engineering principles and general formulas and use.

c.) assuming all energy is converted to heat how much water is needed to absorb the energy lost to heat per day. Assume the water is put back into a stream after heating and the fish in the receiving stream can only withstand temperatures up to 22 degrees

d.) if you were to consider the entire life cycle of producing energy from natural gas, what other flows and process must be considered?
Answer #1

Mass of natural gas used per day = 505 Mg

1 Mg = 1000 Kg

So mass used = 505000 Kg

The energy value of natural gas = 37.3 MJ/kg

Total energy produced = 505000(37.3)= 18836500 MJ= 1.88365 ×1013 J

Electricity produced = 1.81×106 KWh

Electricity produced = 1.81×106(1000) (3600) = 6.516×1012 J

A) Efficiency of plant

= 1-((6.516×1012) /(1.88365×1013) ) = 0.6540 = 65.40%

B) according to lae of conservation of energy

Energy can neither be created nor destroyed hence

Energy lost =

(1.88365×1013) -(6.516×1012) = 1.23205 ×1013 J

C) all lost energy is converted to heat

Q= 1.23205×1013 J

Assume initial temperature of water is 10 °C

Cp = 4.18 KJ/Kg°C

1.23205×1013 = m(4.18×1000)(22-10)

m= 245624.0 Mg

D) if entire life cycle of producing natural gas is considered

The processes that needs to be considered

Purification of natural gas

Compression of natural gas

Combustion of natural gas

Turbine which extracts energy from combustion

All efficiencies of compression, combustion and energy extraction need to be taken care of

Please upvote if helpful